RC meeting and Run Plan, Friday January 20th
Return to the RC Meeting Calendar
RC Meeting Information
|Current RC:||Iuliia Skorodumina|
|Time:||Daily at 3:30PM|
|Location:||Counting House Conference Room|
Remote Connection Details
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
Swing Shift (01/19). We resumed the beam checkout at around 17:30. The harp scans that MCC sent us were great and gave a beam spot on target within our yesterday's specifications (see https://logbooks.jlab.org/entry/4113224). We started the Carbon Hole check, but could not proceed further due to the Moller Target Ion Chamber trips.
MCC and Yves suggested that that our new raster with two sets of coils is causing the trips (see https://logbooks.jlab.org/entry/4113260). They said that the new raster is probably larger than the old raster with one set of coils for which the Ion Chamber Calibration was performed. For this reason, MCC was attempting the new Ion Chamber Calibration for the new raster (see https://logbooks.jlab.org/entry/4113266 and https://logbooks.jlab.org/entry/4113277).
Eventually, it seemed like bleed through from Hall C was likely responsible for the trips (see https://logbooks.jlab.org/entry/4113278), but not the raster, so the beam checkout resumed at around 8pm.
Bill and Jiwan continued the beam checkout until the end of the Swing Shift. They decided to reduce the raster size to achieve better matching with our previous raster with one set of coils (https://logbooks.jlab.org/entry/4113415 and https://logbooks.jlab.org/entry/4113437). At around 1am beam checkout was complete, the new raster size was established and the operational restrictions were updated accordingly. Swing shift summary (https://logbooks.jlab.org/entry/4113171).
Friday Owl (01/20). We started production on Pol He3 Cell with 20 uA beam. At 2:30am MCC went into orbit tuning (with the target at "No target"). At some point beam shifted by 4 mm, which caused BB GEMs HV and Timing Hodoscope HV trips (see https://logbooks.jlab.org/entry/4113540 and https://logbooks.jlab.org/entry/4113551). Sean reset the GEMs HV and tried to reset Hodoscope as well, but the instructions did not seem to be helpful, so they probably need to be updated (https://logbooks.jlab.org/entry/4113550). There is also the issue with a missing BBCal channel (https://logbooks.jlab.org/entry/4113583).
Other than that, everything was great. We took production data at 20 uA, 35 uA, and 45 uA. Owl shift summary (https://logbooks.jlab.org/entry/4113510).
At 7:15 am the beam was taken away as planned and the Target Group decided to do some EPR Calibration before we go into Controlled Access. This seems to be the perfect timing for the EPR Calibration as they do not need beam for that and the Target Polarization can recover during the Day Shift. At some point later today we are going to Controlled Access.
Work in the hall during the Controlled Access, Friday Day Shift (01/20)
- GEMs work. Sean plugged in the cable, so the issue with GEM Layer 2 efficiency is fixed (https://logbooks.jlab.org/entry/4113050).
- HCal work (Jiwan and Sebastian). They were troubleshooting the noisy channels (https://logbooks.jlab.org/entry/4113755) and replacing defective TDC fan (https://logbooks.jlab.org/entry/4113758). They traced the noisy channel issue to originate from the problematic PMT base. They replaced the PMT base, and the noisy channel is now fixed (https://logbooks.jlab.org/entry/4113968).
- Timing Hodoscope (Ralph Marinaro). Fixing the hodoscope issue that occured overnight (see https://logbooks.jlab.org/entry/4113783 and https://logbooks.jlab.org/entry/4113839).
- Checking on gap in BBCal trigger sum (https://logbooks.jlab.org/entry/4113950).
- Work in the Hall during the Controlled Access today (01/20).
- Continue production on pol He3 cell at 45 uA once the beam is back.
- Moller Measurement?
Please consider signing up for shifts .